OpenStudy (lilcathyx):

i just want to make sure that i got it right... Find the sine ratio of angle Θ. Hint: Use the slash symbol ( / ) to represent the fraction bar, and enter the fraction with no spaces.

OpenStudy (lilcathyx):

i think the answer is 15/17

OpenStudy (lilcathyx):

@amistre64 @dan815 am i right ? :o

OpenStudy (johnweldon1993):

The web link doesn't open >.< lol

OpenStudy (lilcathyx):

hold on ;o

OpenStudy (johnweldon1993):

Okay but I'm gonna hold my breath so hurry ;P

OpenStudy (johnweldon1993):

AND YOU AREEEEEE.... Correct :) great job! \(\large \sin(\theta) = \frac{opposite}{hypotenuse}\)

OpenStudy (lilcathyx):

:O Thank you! :D can u check 1 more ? :O

OpenStudy (johnweldon1993):

hmm......maybe...let me think :P haha of course :D

OpenStudy (lilcathyx):

Thank youu ! :D

OpenStudy (lilcathyx):

Find the tangent ratio of angle Θ. Hint: Use the slash symbol ( / ) to represent the fraction bar, and enter the fraction with no spaces.

OpenStudy (lilcathyx):

my cousin says its / but i don't think its that.

OpenStudy (lilcathyx):

12/5 *

OpenStudy (johnweldon1993):

where did the 12 come from? O.o we're using that second triangle right? 

OpenStudy (lilcathyx):

yes that one.

OpenStudy (johnweldon1993):

Alright well I'm not sure where the 12 came from...oh well.. what do YOU think it is? :)

OpenStudy (lilcathyx):

well i think its 4/3..

OpenStudy (johnweldon1993):

That's because you're awesome :) correct indeed ^_^

OpenStudy (johnweldon1993):

\[\large \tan(\theta) = \frac{opposite}{adjacent}\]

OpenStudy (lilcathyx):

:O Omg ! Thank you thank you ! :D

OpenStudy (johnweldon1993):

Anytime hun :) also, there is a quick easy way to remember which is sin, cos and tangent....do you know it?

OpenStudy (lilcathyx):

i dont think so. no i don't

OpenStudy (johnweldon1993):

Just remember SOH CAH TOA S in O pposite H hypotenuse C os A djacent H ypotenuse T an O opposite A adjacent Think of 'S'ome 'O'ld 'H'orse, 'C'aught 'A'nother 'H'orse, 'T'ripping 'O'n 'A'cid

OpenStudy (lilcathyx):

Ohhhh Smarttt :)

OpenStudy (johnweldon1993):

haha I learned that back in like...6th grade i think haha, and I still remember it to this day as a junior in university :D

OpenStudy (lilcathyx):

Omg ! :D thats awesome ! :P Thank you for sharing that with me ! :) ^.^ :D

OpenStudy (johnweldon1993):

Wel of course :D ^_^